OpenStudy (anonymous):

How many solutions does the system have? x = -4y + 4 2x + 8y = 8

OpenStudy (hoblos):

x = -4y + 4 2x + 8y = 8 --> this one can be written as 2x=-8y+8 thus the system is x = -4y + 4 2x = - 8y + 8 notice that if you multiply the first equation by 2 you'll get the second thus it is an infinite system having infinite number of solutions

OpenStudy (anonymous):

Thank that really helped I was stuck